# Scalable Fine-tuning from Multiple Data Sources: A First-Order Approximation Approach

### v2

@@ D.3 Comparison of GPU hours
D.4 Ablation studies
@@
Table 8: Distance vs. speedup for varied $d$, computed using Alpaca and StrategyQA.
@@
Number of subsets $m$:
@@
One might hypothesize that using 1000 or more subsets achieves comparable performance. We conduct the selection algorithm using $m=500,1000,1500,2000$. We notice no further gain in the downstream accuracy when using more subsets than 1000. Thus, we set $m=1000$.
Subset size:
We vary subset size $\alpha$ between $0.25n$, $0.5n$, and $0.75n$. We find the $T_{i}$s all converge at a similar rate. In addition, using a subset size of $0.75n$ achieves better performance in downstream selection. The reason is that this size is closer to the size of the selected subset (i.e., $0.7n$).
D.5 Complete pseudocode for subset selection procedures

### v3

@@ C.1 Runtime-accuracy tradeoff
C.2 Ablation studies of selection algorithms
@@
Table 8: Fine-tuning distance vs. the rate of speedup for various values of dimension $d$, computed using the Alpaca and StrategyQA datasets. For both settings, we find that setting $d=100$ yields the best results. Hence, we use $d=100$ in our experiments for the random projection.
@@
Number of random subsets $m$ in the ensemble method.
@@
One might hypothesize that using 1000 or more subsets achieves comparable performance. To test this, we run the subset selection algorithm using $m=500,1000,1500,2000$. We notice no further performance gain in the downstream test accuracy when using more subsets than $1000$. Thus, we set $m=1000$ in the random ensemble method.
Choice of subset size in random ensemble.
We vary subset size $\alpha$ between $0.25n$, $0.5n$, and $0.75n$. We find the $T_{i}$ scores all converge at a similar rate. Additionally, using a subset size of $0.75n$ yields better performance in downstream selection. The reason is that this size is closer to the size of the selected subset (i.e., $0.7n$).
C.3 Complete subset selection pseudocodes
We provide the complete procedure for forward search and random ensemble with gradient estimation below. In particular, both procedures use Algorithm 1 as a subroutine. For further details, see our code implementation online at https://github.com/VirtuosoResearch/Scalable-finetuning.